OpenStudy (anonymous):

Simplify the expression (5s+st)/(7s)

OpenStudy (lgbasallote):

factor out s from numerator and denominator

OpenStudy (saifoo.khan):

\[\frac{5s +S t }{7s} \to \frac{s(5+t)}{7s}\to \frac{\cancel s(5+t)}{7\cancel s}\to \frac{5+t}{7}\]

OpenStudy (saifoo.khan):

\[\Huge {\cancel{ lgbasallote}}\]@lgbasallote

OpenStudy (lgbasallote):

\(\mathsf{\text{ nice latex}}\)

OpenStudy (saifoo.khan):

\[\huge ikr.\]